OpenStudy (anonymous):

Paraceratherium was the biggest land mammal there has ever been. It lived about 35 million years ago and was 8 m tall and 11 m long. It looked like a giant rhinoceros but had a long neck like a giraffe. About how tall was Paraceratherium in feet? Use the conversion factor 1 ft. = 0.3048 m. Round to the nearest tenth where necessary. a. 26.7 ft. c. 42.6 ft. b. 36.7 ft. d. 19.9 ft. gIVING A MEDAL

OpenStudy (anonymous):

use conversion factors: \[8 m \times \frac{ 1 ft }{ 0.3048m } = \frac{ 8 }{ 0.3048 } ft\] i chose the fraction i did to "cancel out" the units for meters